Find the magnitude of < x2,2> plus <2,x2>
Answer by Edwin McCravy(20054)


< x²,2> + <2,x²> = <2+x²,2+x²>
                 _______________    _______         _
∥<2+x²,2+x²>∥ = Ö(2+x²)²+(2+x²)² = Ö2(2+x)² = |2+x|Ö2

Let's get the dot product of <4,2,3> and <2,-1,-2> 

<4,2,3>•<2,-1,-2> = (4)(2)+(2)(-1)+(3)(-2) = 8-2-6 = 0

So they are orthogonal (perpendicular). 

Edwin